OpenStudy (anonymous):

A ball is dropped from the top of a 1,000 foot building. The height of the ball is half its original height after each bounce. Part 1: What will the height of the ball be after 9 bounces? Part 2: Using complete sentences, explain the procedure taken to answer this question.

OpenStudy (anonymous):

almost same lol as last one.

OpenStudy (anonymous):

I know but I can't plug everything in for some reason

OpenStudy (anonymous):

just remember ar^0 + ar^1 + ar^2 ... 0 1st 2nd ... bounce ^ here we start at 0 instead of 1 cuzz the ball hasn't bounced yet. 1000 +500+250 <- cuzz of the half with me so far?

OpenStudy (anonymous):

69

OpenStudy (anonymous):

tim086m your a virgin aren't you... lol

OpenStudy (anonymous):

512000?

OpenStudy (anonymous):

no i did it wrong let me recheck

OpenStudy (anonymous):

okayy

OpenStudy (anonymous):

the 9th bounce should be ar^9 but before we deal with that lets look at what we do know a=1000 ar^1=ar=500 So we plug in the a and solve for r 1000*r=500 r=500/1000 = 1/2 HERE WHERE I WENT WRONG LAST TIME So we know a=1000 r=1/2 so we get ar^9 and plug in 1000*.5^9= answer :)

OpenStudy (anonymous):

1.953125 ?

OpenStudy (anonymous):

last time i went r=1000/500=2 :P lol it was reverse order. and yes that looks right LB :)

OpenStudy (anonymous):

@blow-some-smoke-420 how you get 69? are you sure? :(

OpenStudy (anonymous):

thankyouuu(:

OpenStudy (anonymous):

im think so why what did u get

OpenStudy (anonymous):

1.953125

OpenStudy (anonymous):

ohh well I think ur right